OpenStudy (anonymous):

The midpoint of A (-4, 2) and B(8, 5) is?

OpenStudy (ivettef365):

use midpoint formula

hartnn (hartnn):

The co-ordinates of midpoint(x,y) on the line segment with end-points (x1,y1) and (x2,y2) is just the average of co-ordinates, given by : \(\large x=\frac{x_1+x_2}{2}\) \(\large y=\frac{y_1+y_2}{2}\) now,just put the values and find the co-ordinates of midpoint (x,y).

OpenStudy (anonymous):

y1-y2 -------- x1-x2?

hartnn (hartnn):

thats slope formula....

OpenStudy (anonymous):

Ughhhh.

OpenStudy (anonymous):

:|

hartnn (hartnn):

for mid-point, just take the average...

hartnn (hartnn):

(-4+8)/2=......?

hartnn (hartnn):

and (5+2)/2 =... ?

OpenStudy (anonymous):

I got 3.5 and 2.

OpenStudy (anonymous):

But the answers to select are (2, 7/2) (-6, 3/2) (4, 7/2)

hartnn (hartnn):

-8+4/2 = 2 5+2/2 =7/2 2,7/2 3.5=7/2

OpenStudy (anonymous):

So would I round to (4,7/2)?

hartnn (hartnn):

(-4+8)/2 = 4/2 = 2 so it would be (2,7/2)

OpenStudy (anonymous):

Ohhhhh.Thank you!

hartnn (hartnn):

welcome ^_^